OpenStudy (routohan75):

I AM SO BORED SOME ONE PLEASE HELP ME WITH MY MATH!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

Yes I can @Routohan75

OpenStudy (routohan75):

ok

OpenStudy (routohan75):

What is the point-slope form of the equation of a line that passes through the point (1, 4) and has a slope of −3? A. y-1= -3(x-4) B. y+4=3(x+1) C. y+4= -3(x+1) D. y-4 =-3(x-1)

OpenStudy (anonymous):

What would ur answer be @Routohan75

OpenStudy (routohan75):

b or c but I suck with math so I know I am going to be wrong

OpenStudy (anonymous):

C is the correct answer, if you would have picked c as ur lucky letter you would have been correct.

OpenStudy (routohan75):

THANK YOU!!

Nnesha (nnesha):

well bec C for chocolates that's why lucky ??

OpenStudy (anonymous):

You are very welcome @Routohan75 anytime

OpenStudy (anonymous):

Wow lol @Nnesha

Nnesha (nnesha):

\[\huge\rm y-y_1 =m(x-x_1)\] point slope form where m is slope and (x_1 ,y_1) is the point

OpenStudy (anonymous):

Chocolate of course

Nnesha (nnesha):

sorry to say C is wrong.

OpenStudy (anonymous):

How come?

Nnesha (nnesha):

look at the original equation for point slope plug in -3 for m and passes through the point (1,4) for (x_1 ,y_1)

OpenStudy (anonymous):

Sorry gotta go but thanks for correcting me @Nnesha because I really do not like giving people the wrong answer.

OpenStudy (routohan75):

it is not c?

Nnesha (nnesha):

np

OpenStudy (routohan75):

What would it be then?

Nnesha (nnesha):

i gave you the original formula for point slope just plugin numbers and see which one would be the correct one